A tank is filled with water to a height of 12.5 cm. The apparent depth of a needle lying at the bottom of the tank is measured by a microscope to be 9.4 cm. What is the refractive index of water? If water is replaced by a liquid of refractive index 1.63 up to the same height, by what distance would the microscope have to be moved to focus on the needle again?

Grade:11

1 Answers

Vikas TU
3 years ago
Apparent depth = real depth / refractive index
or
9.4 = 12.5 / μ
μ = 12.5 /9.4
μ = 1.33.

Now for the other part.
Apparent depth h' = 12.5 / μ'
h' = 12.5 /1.63
h' = 7.67

Previously apparent depth was 9.4 cm
Hence Microscope have to be moved by
= (9.4 - 7.67 )
= 1.73 cm.